Sexton Inc. is considering Projects S and L, whose cash flows are shown below. These projects are mutually exclusive, equally risky, and not repeatable. If the decision is made by choosing the project with the higher IRR, how much value will be forgone? Note that under certain conditions choosing projects on the basis of the IRR will not cause any value to be lost because the one with the higher IRR will also have the higher NPV, so no value will be lost if the IRR method is used. Sexton Inc. is considering Projects S and L, whose cash flows are shown below. These projects are mutually exclusive, equally risky, and not repeatable. If the decision is made by choosing the project with the higher IRR, how much value will be forgone? Note that under certain conditions choosing projects on the basis of the IRR will not cause any value to be lost because the one with the higher IRR will also have the higher NPV, so no value will be lost if the IRR method is used WACC: 9.50% 4 CFs $2,050 $750 $70 $770 $780 CFL -$4,300 $1,500 $1,518$1,536 $1,554 $188.91 $145.46 S228.58 $226.70 $230.47
Answer #1

We need to find the NPV of both projects and find the difference:

NPV of project S = Present value of cash inflows - present value of cash outflows

NPV of project S = -2,050 + 750 / ( 1 + 0.095)1 + 760 / ( 1 + 0.095)2 + 770 / ( 1 + 0.095)3 + 780 / ( 1 + 0.095)4

NPV of project S = $397.801

NPV of project L = Present value of cash inflows - present value of cash outflows

NPV of project L = -4,300 + 1,500 / ( 1 + 0.095)1 + 1,518 / ( 1 + 0.095)2 + 1,536 / ( 1 + 0.095)3 + 1,554 / ( 1 + 0.095)4

NPV of project L = $586.714

Lost value = 586.714 - 397.801 = $188.91